For a segment of a radio show, a disc jockey can play 6 records.If there are 11 records to choose from in how many ways can a segment for this program be arranged?
Answer by stanbon(75887)

For a segment of a radio show, a disc jockey can play 6 records.If there are 11 records to choose from in how many ways can a segment for this program be arranged?
------
11P6 = 11!/(11-6)! = 11!/5! = 11*10..*6 = 332640
----------
Cheers,
Stan H.
--------